Grade 9Mechanics

A ball thrown vertically upwards is caught in time t second.find height

S N TRIPATHI
7 Years agoGrade 9

2 Answers

Arun
7 Years ago
Dear student
 
as time of flight when going upward = t/2
hence
 
h = 0* t + ½ g (t/2)^2
 
h = 5t2/4
 
Hope it helps
Khimraj
7 Years ago
t/2 sec for the first path meaning when up until the ball reaches the max height.
Now, h=ut+ 1/2*g*(t/2)^2
u = 0, g=10
so h = 5t2/4